Joe traveled against the wind in a small plane for 3 hours. The return trip with the wind took 2.8 hours. What is the speed of the wind if the speed of the plane in still air is 180 mph?
Answer by mananth(16949)

Joe traveled against the wind in a small plane for 3 hours. The return trip with the wind took 2.8 hours. What is the speed of the wind if the speed of the plane in still air is 180 mph?
Let the speed of the wind be x
Against the wind the speed will be 180-x
Distance traveled will be time * speed
=3*(180-x)
With wind the speed will be 180 +x
Distance traveled will be = 2.8* (180+x)
The distance traveled will be same
2.8*(180+x)= 3*(180-x)
504+2.8x=540-3x
2.8x+3x=540-504
5.8x=36
x= 6.2 mph is the speed of the wind.